how do I solve 4(2n-5)=3n+10
Found 3 solutions by erica65404, Alan3354, angel.wit.de.halo:
Answer by erica65404(394)
Answer by Alan3354(69443)

solve 4(2n-5)=3n+10
8n - 20 = 3n + 10
8n = 3n + 30
5n = 30
n = 6

Answer by angel.wit.de.halo(6)

8n-20=3n+10...add twenty to both sides to get the numbers to one side
8n=3n+30...subtract 3n from both sides to get the variables to one side
5n=30...divide 5 on both sides to get n=...
n=6